# Progressive Collaborative and Semantic Knowledge Fusion for Generative Recommendation

### v1

@@ Abstract.
With the recent surge in interest surrounding generative paradigms, generative recommendation has increasingly attracted the attention of researchers in the recommendation community. This paradigm generally consists of two stages. In the first stage, pretrained semantic embeddings or collaborative ID embeddings are quantized to create item codes, aiming to capture and preserve rich semantic or collaborative knowledge within these codes. The second stage involves utilizing these discrete codes to perform an autoregressive sequence generation task. Existing generative recommendation methods often either overlook collaborative knowledge or semantic knowledge, or combine the two roughly. In this paper, we observe that naively concatenating representations from semantic modality and collaborative modality leads to a semantic domination issue, where the resulting representation is overly influenced by semantic information, effectively overshadowing the collaborative representation. Consequently, downstream recommendation tasks fail to fully exploit the knowledge from both modalities, resulting in suboptimal performance.
To address this, we propose a progressive collaborative and semantic knowledge fusion model for generative recommendation, named PRORec, which integrates semantic and collaborative knowledge with a unified code through a two-stage framework. Specifically, in the first stage, we propose a cross-modality knowledge alignment task during the pre-training of the non-generative recommendation model, which integrates semantic knowledge into collaborative embeddings, enhancing their representational capability. In the second stage, we propose an in-modality knowledge distillation task, designed to encourage the generative recommendation model to effectively capture and integrate knowledge from both semantic and collaborative modalities. Extensive experiments on three widely used benchmarks validate the effectiveness of our approach, demonstrating its superiority compared to existing methods.
@@ 1. INTRODUCTION
To overcome the aforementioned challenges, we propose a novel method, PRORec, which integrates collaborative and semantic knowledge into a unified code for generative recommendation. Unified codes not only reduce storage and computation demands but also accelerate inference and lower latency, which are critical for real-time responses. As illustrated in Figure 3, a unified code demonstrates significantly 2.8x faster inference compared to two separate codes. Moreover, dual-code methods inherently treat semantic and collaborative knowledge as independent entities, thereby missing opportunities to harness their synergistic potential. In contrast, a unified code, by integrating both knowledge at an early stage, can more effectively exploit their complementary strengths.
@@
  • We introduce PRORec, with a unified code, Unicodes, which integrates both collaborative and semantic knowledge for generative recommendation. By utilizing the unified code, we achieve faster inference while requiring only half the storage space compared to two separate codes setup.
@@
  • Extensive experiments on three public recommendation benchmarks demonstrate PRORec’s superiority over existing methods, encompassing both generative and traditional paradigms. Additionally, we observe that PRORec demonstrates the scaling law characteristic.
@@ 3.2. Overall Pipeline
We present PRORec overall framework in Figure 4, which consists of two stages. In the first stage, we first adaptively learn integrated embeddings through the joint optimization of proposed cross-modality knowledge alignment task and next item prediction task, which fuse knowledge from both collaborative and semantic modalities. Subsequently, the integrated embeddings are quantized into discrete unicodes that act as item identifiers in the second stage. The second stage is the generative recommendation phase, comprising an encoder, a decoder and an intra-modality knowledge distillation task to compensate for the information loss caused by the quantization process. Here, the user’s historical interaction sequence is transformed into unicodes sequence using the item-unicode mapping table obtained in the first stage. The encoder models the user’s historical interactions to capture their interests, which are then passed into the decoder to guide the generation of unicodes for candidate items. Once training is complete, we select the top-$k$ recommendations based on confidence scores.
Figure 4. An overview of PRORec. PRORec consists of two stages. The first stage integrates semantic and collaborative knowledge to construct the unified code, Unicode for each item. The second stage utilizes the obtained unicodes to perform generative recommendation. To achieve the goal of utilizing a unified code to encode the two different knowledge for generative recommendation, we introduce two auxiliary tasks at each stage: a cross-modality knowledge alignment task (CKA) in the first stage and an intra-modality knowledge distillation task (IKD) in the second stage. Besides, in the first stage, a modality adaption layer with AdaLN is also introduced to bridge the modality gap between semantic and collaborative space.
@@ 3.6. Computational and Storage Costs Analysis
TIGER EAGER PRORec (Ours)
@@
Table 3. Performance comparison of different methods. The best performance is highlighted in bold while the second best performance is underlined. All the results of PRORec are statistically significant with $p<0.05$ compared to the best baseline models.
Dataset Metric Classical Generative PRORec (Ours)
GRU4REC Caser SASRec BERT4Rec HGN FDSA S3-Rec Recorest TIGER EAGER
Beauty Recall@10 0.0283 0.0347 0.0605 0.0347 0.0512 0.0407 0.0647 0.0664 $0.0617^{*}$ 0.0836 0.0939
Recall@20 0.0479 0.0556 0.0902 0.0599 0.0773 0.0656 0.0994 0.0915 $0.0924^{*}$ 0.1124 0.1289
NDCG@10 0.0137 0.0176 0.0318 0.0170 0.0266 0.0208 0.0327 0.0400 $0.0339^{*}$ 0.0525 0.0559
NDCG@20 0.0187 0.0229 0.0394 0.0233 0.0332 0.0270 0.0414 0.0464 $0.0417^{*}$ 0.0599 0.0646
Sports Recall@10 0.0204 0.0194 0.0350 0.0191 0.0313 0.0288 0.0385 0.0247 $0.0376^{*}$ 0.0441 0.0471
Recall@20 0.0333 0.0314 0.0507 0.0315 0.0477 0.0463 0.0607 0.0375 $0.0577^{*}$ 0.0659 0.0710
NDCG@10 0.0110 0.0097 0.0192 0.0099 0.0159 0.0156 0.0204 0.0133 $0.0196^{*}$ 0.0236 0.0259
NDCG@20 0.0142 0.0126 0.0231 0.0130 0.0201 0.0200 0.0260 0.0164 $0.0246^{*}$ 0.0291 0.0319
Toys Recall@10 0.0176 0.0270 0.0675 0.0203 0.0497 0.0381 0.0700 0.0383 $0.0578^{*}$ 0.0714 0.0822
Recall@20 0.0301 0.0420 0.0941 0.0358 0.0716 0.0632 0.1065 0.0483 $0.0838^{*}$ 0.1024 0.1154
NDCG@10 0.0084 0.0141 0.0374 0.0099 0.0277 0.0189 0.0376 0.0285 $0.0321^{*}$ 0.0505 0.0489
NDCG@20 0.0116 0.0179 0.0441 0.0138 0.0332 0.0252 0.0468 0.0310 $0.0386^{*}$ 0.0538 0.0573
@@ 4. EXPERIMENTS
We analyze the proposed PRORec method and demonstrate its effectiveness by answering the following research questions:
  • RQ1: How does PRORec perform compared with existing best-performing sequential recommendation methods among different datasets?
  • RQ2: How effective is PRORec in mitigating the semantic domination issue?
  • RQ3: Do cross-modality knowledge alignment and intra-modality knowledge distillation tasks each contribute positively to PRORec’s performance?
  • RQ4: How do different ablation variants and hyper-parameter settings affect the performance of PRORec?
@@ 4.1. Experimental Setting
Implementation Details. In our experiments, we set the number of encoder layers to 1 and the number of decoder layers to 4. Following previous work (Feng et al., 2022; Wang et al., 2024b), we adopt DIN as our sequential recommendation model and use pretrained Llama2-7b as our semantic encoder, with the hidden size set to 128 as reported in (Rajput et al., 2023; Wang et al., 2024b). The number of clusters in the hierarchical k-means is set to 256. To train our model, we use the Adam optimizer with a learning rate of 0.001 and apply a warmup strategy for stable training. The batch size is set to 256. Notably, PRORec demonstrates robustness to the hyperparameters of the CKA and IKD tasks, attributed to their fast convergence. Consequently, the loss coefficients $\alpha$ and $\beta$ are both set to 1.
@@ 4.2. Performance Comparison (RQ1)
Baselines.We compare PRORec with several representative related baselines, encompassing classical sequential modeling approaches as well as the latest emergent generative techniques.
@@
Results. Table 3 reports the overall performance of three datasets. The results for all baselines without the superscript are taken from the publicly accessible results (Zhou et al., 2020; Wang et al., 2024b). Due to missing values for TIGER, we reimplement it and report our experimental results. From the results, we have the following observations:
  • PRORec almost achieves better results than base models among different datasets. Specifically, on the Beauty dataset, compared to the second best baseline EAGER, we achieve improvements of 14.68% and 7.85% in Recall@20 and NDCG@20, respectively, which represents a significant improvement in the recommendation field. We attribute the improvements to PRORec’s effective exploitation of the complementary strengths of collaborative and semantic knowledge with a unified code, which allows it to achieve superior performance relative to EAGER, despite utilizing approximately half the number of total parameters.
  • PRORec beats the previous generative models on most datasets. Compared to previous generative models, PRORec employs a two-stage multi-task joint optimization framework to integrate collaborative and semantic knowledge with a unified code, which successfully captures the interplay between these two types of knowledge, allowing for a more thorough and efficient exploitation of their complementary strengths. The improvements validate the effectiveness of our method and demonstrate the necessity of a unified code.
@@
Semantic Only 0.0810 0.1102 0.0498 0.0571
Collaborative Only 0.0803 0.1095 0.0495 0.0569
Concat 0.0807 0.1085 0.0504 0.0574
@@ 4.3. Semantic Domination Issue (RQ2)
Beauty Recall@10 0.0807 0.0827 0.0939
Recall@20 0.1085 0.1122 0.1289
NDCG@10 0.0504 0.0509 0.0559
NDCG@20 0.0574 0.0583 0.0646
@@ 4.4. Ablation Study (RQ3)
We conducted a thorough evaluation of the contribution of each component in PRORec. Specifically, we incrementally removed the intra-modality knowledge distillation (IKD) task and the cross-modality knowledge alignment (CKA) task to create the ablation architectures. The results in Table 6 reveal the following insights:
@@
In conclusion, each component of PRORec is essential to enhancing the effectiveness of generative recommendation.
@@ 5. CONCLUSION AND FUTURE WORK
In this paper, we introduce a novel framework, PRORec, aimed at leveraging a unified code to encode knowledge from both semantic and collaborative modalities in generative recommendation systems. PRORec is composed of three key components: (1) A unified code generation framework that integrates semantic and collaborative knowledge, addressing the semantic domination issue by effectively capturing their complementary characteristics. (2) A joint optimization of cross-modality knowledge alignment and next item prediction tasks and a learnable modality adaption layer with AdaLN, which integrate collaborative and semantic knowledge into integrated embeddings, enabling the unified code to represent both modalities information cohesively. (3) An intra-modality knowledge distillation task with a specially designed token, which compensates for the information loss caused by the quantization process and further improves auto-regressive generation quality. Extensive experiments compared with state-of-the-art methods, combined with detailed analyses, validate the effectiveness and robustness of PRORec. In future work, we plan to leverage more available features from the semantic modality, such as item visual features, to further enhance the model’s generative recommendation performance.

### v2

# UNGER: Generative Recommendation with A Unified Code via Semantic and Collaborative Integration

@@ Abstract.
With the recent surge in the popularity of generative paradigms, generative recommendation has gradually drawn attention from researchers in the recommendation community. The core component in generative recommendation is the item code, which is generally acquired through the quantization of collaborative or semantic modality representation, and then acts as the candidate item identifier in the auto-regressive generative model for sequence-to-sequence modeling. However, existing methods typically construct independent codes for different modalities, which not only double the computational and storage costs but also impede the effective integration of the complementary strengths of collaborative and semantic knowledge. Considering this limitation, this paper seeks to integrate two different modalities into a unified code, fully unleashing the potential of complementary nature among modalities. Nevertheless, the integration remains challenging: the integrated embedding obtained by the common concatenation method aligns predominantly with semantic representation, overshadowing collaborative information. This imbalance would lead to underutilization of collaborative knowledge, thereby resulting in limited effectiveness.
To address this, we propose a novel method, named UNGER, which integrates semantic and collaborative knowledge into a unified code for generative recommendation. Specifically, we propose to adaptively learn an integrated embedding through the joint optimization of cross-modality knowledge alignment and next item prediction tasks. Subsequently, to mitigate the information loss caused by the quantization process, we further introduce an intra-modality knowledge distillation task, using the integrated embeddings as supervised signals to compensate during the training of the auto-regressive generative model. Extensive experiments on three widely used benchmarks demonstrate the superiority of our approach compared to existing methods, including both generative and traditional paradigms.
@@ 1. INTRODUCTION
To overcome the aforementioned challenges, we propose a novel method, UNGER, which integrates collaborative and semantic knowledge into a unified code for generative recommendation. Unified codes not only reduce storage and computation demands but also accelerate inference and lower latency, which are critical for real-time responses. As illustrated in Figure 3, a unified code demonstrates significantly 2.8x faster inference compared to two separate codes. Moreover, dual-code methods inherently treat semantic and collaborative knowledge as independent entities, thereby missing opportunities to harness their synergistic potential. In contrast, a unified code, by integrating both knowledge at an early stage, can more effectively exploit their complementary strengths.
@@
  • We introduce UNGER, with a unified code, Unicodes, which integrates both collaborative and semantic knowledge for generative recommendation. By utilizing the unified code, we achieve faster inference while requiring only half the storage space compared to two separate codes setup.
@@
  • Extensive experiments on three public recommendation benchmarks demonstrate UNGER’s superiority over existing methods, encompassing both generative and traditional paradigms. Additionally, we observe that UNGER demonstrates the scaling law characteristic.
@@ 3.2. Overall Pipeline
We present UNGER overall framework in Figure 4, which consists of two stages. In the first stage, we first adaptively learn integrated embeddings through the joint optimization of proposed cross-modality knowledge alignment task and next item prediction task, which fuse knowledge from both collaborative and semantic modalities. Subsequently, the integrated embeddings are quantized into discrete unicodes that act as item identifiers in the second stage. The second stage is the generative recommendation phase, comprising an encoder, a decoder and an intra-modality knowledge distillation task to compensate for the information loss caused by the quantization process. Here, the user’s historical interaction sequence is transformed into unicodes sequence using the item-unicode mapping table obtained in the first stage. The encoder models the user’s historical interactions to capture their interests, which are then passed into the decoder to guide the generation of unicodes for candidate items. Once training is complete, we select the top-$k$ recommendations based on confidence scores.
Figure 4. An overview of UNGER. UNGER consists of two stages. The first stage integrates semantic and collaborative knowledge to construct the unified code, Unicode for each item. The second stage utilizes the obtained unicodes to perform generative recommendation. To achieve the goal of utilizing a unified code to encode the two different knowledge for generative recommendation, we introduce two auxiliary tasks at each stage: a cross-modality knowledge alignment task (CKA) in the first stage and an intra-modality knowledge distillation task (IKD) in the second stage. Besides, in the first stage, a modality adaption layer with AdaLN is also introduced to bridge the modality gap between semantic and collaborative space.
@@ 3.6. Computational and Storage Costs Analysis
TIGER EAGER UNGER (Ours)
@@
Table 3. Performance comparison of different methods. The best performance is highlighted in bold while the second best performance is underlined. All the results of UNGER are statistically significant with $p<0.05$ compared to the best baseline models.
Dataset Metric Classical Generative UNGER (Ours)
GRU4REC Caser SASRec BERT4Rec HGN FDSA S3-Rec Recorest TIGER ColaRec EAGER
Beauty Recall@10 0.0283 0.0347 0.0605 0.0347 0.0512 0.0407 0.0647 0.0664 $0.0617^{*}$ $0.0524^{*}$ 0.0836 0.0939
Recall@20 0.0479 0.0556 0.0902 0.0599 0.0773 0.0656 0.0994 0.0915 $0.0924^{*}$ $0.0807^{*}$ 0.1124 0.1289
NDCG@10 0.0137 0.0176 0.0318 0.0170 0.0266 0.0208 0.0327 0.0400 $0.0339^{*}$ $0.0263^{*}$ 0.0525 0.0559
NDCG@20 0.0187 0.0229 0.0394 0.0233 0.0332 0.0270 0.0414 0.0464 $0.0417^{*}$ $0.0335^{*}$ 0.0599 0.0646
Sports Recall@10 0.0204 0.0194 0.0350 0.0191 0.0313 0.0288 0.0385 0.0247 $0.0376^{*}$ $0.0348^{*}$ 0.0441 0.0471
Recall@20 0.0333 0.0314 0.0507 0.0315 0.0477 0.0463 0.0607 0.0375 $0.0577^{*}$ $0.0533^{*}$ 0.0659 0.0710
NDCG@10 0.0110 0.0097 0.0192 0.0099 0.0159 0.0156 0.0204 0.0133 $0.0196^{*}$ $0.0179^{*}$ 0.0236 0.0259
NDCG@20 0.0142 0.0126 0.0231 0.0130 0.0201 0.0200 0.0260 0.0164 $0.0246^{*}$ $0.0226^{*}$ 0.0291 0.0319
Toys Recall@10 0.0176 0.0270 0.0675 0.0203 0.0497 0.0381 0.0700 0.0383 $0.0578^{*}$ $0.0474^{*}$ 0.0714 0.0822
Recall@20 0.0301 0.0420 0.0941 0.0358 0.0716 0.0632 0.1065 0.0483 $0.0838^{*}$ $0.0704^{*}$ 0.1024 0.1154
NDCG@10 0.0084 0.0141 0.0374 0.0099 0.0277 0.0189 0.0376 0.0285 $0.0321^{*}$ $0.0242^{*}$ 0.0505 0.0489
NDCG@20 0.0116 0.0179 0.0441 0.0138 0.0332 0.0252 0.0468 0.0310 $0.0386^{*}$ $0.0300^{*}$ 0.0538 0.0573
@@ 4. EXPERIMENTS
We analyze the proposed UNGER method and demonstrate its effectiveness by answering the following research questions:
  • RQ1: How does UNGER perform compared with existing best-performing sequential recommendation methods among different datasets?
  • RQ2: How effective is UNGER in mitigating the semantic domination issue?
  • RQ3: Do cross-modality knowledge alignment and intra-modality knowledge distillation tasks each contribute positively to UNGER’s performance?
  • RQ4: How do different ablation variants and hyper-parameter settings affect the performance of UNGER?
@@ 4.1. Experimental Setting
Implementation Details. In our experiments, we set the number of encoder layers to 1 and the number of decoder layers to 4. Following previous work (Feng et al., 2022; Wang et al., 2024b), we adopt DIN as our sequential recommendation model and use pretrained Llama2-7b as our semantic encoder, with the hidden size set to 128 as reported in (Rajput et al., 2023; Wang et al., 2024b). The number of clusters in the hierarchical k-means is set to 256. To train our model, we use the Adam optimizer with a learning rate of 0.001 and apply a warmup strategy for stable training. The batch size is set to 256. Notably, UNGER demonstrates robustness to the hyperparameters of the CKA and IKD tasks, attributed to their fast convergence. Consequently, the loss coefficients $\alpha$ and $\beta$ are both set to 1.
@@ 4.2. Performance Comparison (RQ1)
Baselines.We compare UNGER with several representative related baselines, encompassing classical sequential modeling approaches as well as the latest emergent generative techniques.
@@
  • ColaRec (Wang et al., 2024a) integrates user-item interactions and content data within an end-to-end framework, leveraging pretrained collaborative identifiers, an item indexing task, and contrastive loss to align semantic and collaborative spaces for enhanced recommendation performance.
@@
Results. Table 3 reports the overall performance of three datasets. The results for all baselines without the superscript are taken from the publicly accessible results (Zhou et al., 2020; Wang et al., 2024b). For missing statistics, we reimplement the baseline and report our experimental results. From the results, we have the following observations:
  • UNGER almost achieves better results than base models among different datasets. Specifically, on the Beauty dataset, compared to the second best baseline EAGER, we achieve improvements of 14.68% and 7.85% in Recall@20 and NDCG@20, respectively, which represents a significant improvement in the recommendation field. We attribute the improvements to UNGER’s effective exploitation of the complementary strengths of collaborative and semantic knowledge with a unified code, which allows it to achieve superior performance relative to EAGER, despite utilizing approximately half the number of total parameters.
  • UNGER beats the previous generative models on most datasets. Compared to previous generative models, UNGER employs a two-stage multi-task joint optimization framework to integrate collaborative and semantic knowledge with a unified code, which successfully captures the interplay between these two types of knowledge, allowing for a more thorough and efficient exploitation of their complementary strengths. The improvements validate the effectiveness of our method and demonstrate the necessity of a unified code.
@@
Semantic Only 0.0791 0.1077 0.0462 0.0535
Collaborative Only 0.0744 0.0997 0.0469 0.0532
Concat 0.0759 0.1071 0.0452 0.0530
@@ 4.3. Semantic Domination Issue (RQ2)
Beauty Recall@10 0.0759 0.0827 0.0939
Recall@20 0.1071 0.1122 0.1289
NDCG@10 0.0452 0.0509 0.0559
NDCG@20 0.0530 0.0583 0.0646
@@ 4.4. Ablation Study (RQ3)
We conducted a thorough evaluation of the contribution of each component in UNGER. Specifically, we incrementally removed the intra-modality knowledge distillation (IKD) task and the cross-modality knowledge alignment (CKA) task to create the ablation architectures. The results in Table 6 reveal the following insights:
@@
In conclusion, each component of UNGER is essential to enhancing the effectiveness of generative recommendation.
@@ 5. CONCLUSION AND FUTURE WORK
In this paper, we introduce a novel framework, UNGER, aimed at leveraging a unified code to encode knowledge from both semantic and collaborative modalities in generative recommendation systems. UNGER is composed of three key components: (1) A unified code generation framework that integrates semantic and collaborative knowledge, addressing the semantic domination issue by effectively capturing their complementary characteristics. (2) A joint optimization of cross-modality knowledge alignment and next item prediction tasks and a learnable modality adaption layer with AdaLN, which integrate collaborative and semantic knowledge into integrated embeddings, enabling the unified code to represent both modalities information cohesively. (3) An intra-modality knowledge distillation task with a specially designed token, which compensates for the information loss caused by the quantization process and further improves auto-regressive generation quality. Extensive experiments compared with state-of-the-art methods, combined with detailed analyses, validate the effectiveness and robustness of UNGER. In future work, we plan to leverage more available features from the semantic modality, such as item visual features, to further enhance the model’s generative recommendation performance.